# What You Say = What You Want? Teaching Humans to Articulate Requirements for LLMs

### v1

@@ 5. User Study Design
We employed both within-subject and between-subject designs in our study. To capture the requirement-focused training gains, we utilized a pre-test and post-test approach, a standard instruction evaluation method (Rogaten et al., 2019), comparing participants’ performance on isomorphic tasks before and after training. To compare requirement-focused training with standard training, we utilized a randomized-control design. We randomly assigned participants to either the experimental condition (ROPE group), where they receive requirement-focused training in our interface, or the control condition (PE group), where they receive a standard prompt engineering tutorial and then self-practice using ChatGPT.
Our user study — for either the ROPE or the PE group — is 1.5 hours long. In the study, participants first completed a two-minute pre-survey, providing demographic information and rating their prior prompting experiences with 7-level Likert Scale questions. Then, they started a pre-test (20 minutes) to write prompts to replicate key features on existing programs (Table 1) — one Game (Tic-Tac-Toe or Connect4) and one GPTs (Outline Assistant or Trip Advisor). They proceeded to the 40-minute training session (different interfaces depending on experiment group) on two tasks (Tetris and Email Proofreader), followed by a post-test that contained different Game and GPTs tasks (20 minutes). To further capture participants’ iterative prompting behaviors, for the post-test GPTs task, participants were also asked to feed their prompt into ChatGPT and iterate their original prompt by observing the ChatGPT outputs. They completed the study with another two-minute post-survey, providing feedback with open-ended or Likert Scale questions on their experience and perceptions during the training. Participants were compensated with a \$20 Gift Card.
@@
We collected participants’ responses (prompts), time on task, as well as interaction log data during the training phase. Afterward, we collected the LLM outputs by feeding participant prompts into OpenAI GPT-4o. [^10] We used default settings such as a temperature of 1, please refer to our supplemental materials for more details. Using the evaluation method and rubrics we developed for each of the assessment tasks Section 4.2, we graded the pre- and post- tests using users’ original prompts for all tasks from the test, on a scale of 0% to 100%. We calculated the Overall Test Scores as the average of the Requirement Quality and LLM Output Quality scores. The pre- and post- assessments each contained one Game task and one GPTs task (Section 4.1), counterbalanced to mitigate problem sequencing bias or different task difficulties. We found that the two versions of the tests have comparable difficulties from pre-test results (an average of 22.9% and 18.3% for each version).
To understand the effect of an optimizer on prompt (as discussed in Section 2.2), we further use the Prompt Maker, [^11] https://chatgpt.com/g/g-hhh4w3eov-prompt-maker: from a simple prompt to an optimized prompt. a popular optimizer with 50k usages in GPT Store, to refine participants’ original prompts into optimized prompts. This optimizer is easily applicable to any kinds of prompts (Figure 2 shows an example of its output), though it is a rather preliminary version, as it rephrases user prompts in a rule-based manner without criticizing the LLM outputs of the original prompt. We evaluated the optimized prompts in the same way as the human prompts (Section 4.2), including generating and grading corresponding LLM outputs.
@@
We recruited users with limited or no NLP background from different institutions in the United States. We conducted studies with 32 participants randomly assigned to the ROPE or PE group, and we removed two participants who disengaged during the study from the analysis. In total, we have 30 participants (S1-30, $n=15$ for each condition) — 19 female, 10 male, 1 agender, and 18 non-native English speakers, with an average age of 26.
@@ 6.1. Learning Gain: Requirement-focused Training Helps Students Instruct LLMs
Using a two-tailed paired t-test, [^12] We did a linear mixed effects regression to account for the correlation across participants and task (both as random effects) of the repeated measures for the pre- and post-test times and for the requirement and LLM output scores. We found the same pattern of significant results as revealed by the t-tests. we found that test scores for participants in the ROPE group significantly improved by 19.8% from pre- to post-test ($p<0.001$, from $20.0\%\pm 14.6\%$ to $39.8\%\pm 18.8\%$), a two-fold increase. This demonstrated that requirement-focused training enabled participants to instruct LLMs more effectively, and we achieved our desired goal: to concentrate users’ attention on only requirement iterations during the training, but make sure users can still apply their learning in overall prompt writing. Analyzing requirement defects in ROPE participants’ prompts, we observed a noticeable decrease in omission errors (from 5.6 to 3.2 per participant), but a slight increase in commission errors (from 0.5 to 0.7 per participant). This indicated that participants became more aware of requirements, but need additional training for requirements clarity and accuracy.
ROPE participants’ self-reflection highlighted that they understand the value of emphasizing and iterating on requirements, which may contribute to their improvement. Six ROPE participants explicitly noted the connection between requirements and prompts in their post-survey. For example, S30 described requirements as a way “to be more clear and not confuse the system by giving too much unnecessary information”, and S8 noted that a requirement-focused approach helped them organize their thoughts: “I learned to organize my requirements logically so we can easily revise and improve them”. S10 neatly described their shift towards requirement-focused prompting strategy: “sometimes when I write prompts, [I find] steps are hard to be clearly divided, or I didn’t consider to divide them that detailed. However, it’s important to do so, as it appears to give LLM more direct and clear instruction. When the steps are divided, it’s easier to see the missing details in my original prompts too”.
@@
These prompt also revealed learning gain differences. A two-sample t-test showed that the ROPE achieved significantly higher learning gains compared to the PE group ($p<0.001$). In fact, while ROPE significantly improved (19.8% as mentioned above), we did not observe much gain in PE group’s pre-to-post score (1.2%). The contrast suggested that novices indeed could not acquire requirement articulation skills through standard prompting training alone.
@@
Requirement articulation was challenging for all participants, as shown by the low pre-test scores (20% for the ROPE and 23% for the PE). This poor starting point likely hindered the PE’s ability to receive useful feedback during self-practice, as their requirements were too weak for the LLM to generate “good enough” model output with clear indicators on what to improve. Consequently, PE participants were trapped and frustrated by LLM unpredictability — S6 noted that “ChatGPT is a very malleable tool and can change responses pretty drastically depending on the prompt”, and S7 commented that “if ChatGPT doesn’t get it, I might not be patient”.
In contrast, our feedback loop for ROPE (Section 4.3) provided tightly controlled feedback and program output based on the requirements users specified, regardless of grammatical issues or natural language variances. S8 highlighted that: “It’s easy to understand (very logical) and can see how the changes and revises influences the system immediately”. This deterministic feedback reinforced ROPE participants’ believes on the importance of clear requirements, leading to them writing better initial prompts with more requirements in the post-test. We hypothesize that these better initial prompts might also become more suitable for further iteration. Although still imperfect, ROPE participants’ prompts might allow the LLM to generate partially correct outputs that contain enough signal towards future improvements.
@@ 6.2. In-depth Analysis: The Validity of ROPE Paradigm
Figure 5. (A) Spearman correlation ($\rho$) between LLM Output and Requirement Score by Tasks. (B) A comparison between the prompt format sensitivity of Connect4 and Tic-Tac-Toe. We suspect the low correlation Connect4 on requirement vs. LLM output quality is due to the lack of natural-language-to-code mapping in LLM training on this task, making it more sensitive to prompt formats.
[…]
Going beyond learning gains, we dive deeper into the connections between requirement in prompts and LLM output quality, as well as the role of optimizers.
@@
To examine whether more correct and complete requirements led to better LLM outputs, we calculated the correlation between the two components of our Overall Test Scores: Requirement Quality Score and LLM Output Quality Score. We found a strong positive correlation, with a Spearman’s correlation coefficient of $\rho=0.72$.
However, task-specific analysis revealed nuances. As shown in Figure 5, while all other tasks achieved $\rho\geq 0.88$, Connect4 did not show any correlation. Besides one ROPE participant who scored 80% on requirement quality and 70% on LLM output in their post-test Connect4, all other participants’ LLM output scores saturated around 40% even with high requirement quality $\geq$ 80%.
Upon reviewing user prompts and LLM outputs, we suspected that Connect4 had fewer natural-language-to-code mappings in GPT-4o’s training data. [^13] Confirming these hypotheses would need more systematic probing and quantification (Qian et al., 2023), which we defer to future work. Consequently, the model encountered difficulties implementing uncommon requirements in code for Connect4, such as ‘‘Cross out the winning cells’’, which only succeeded in 1 out of 16 prompts. Moreover, as shown in Figure 5, Connect4 prompts formatted as code implementation plans produced better results than casual narrative prompts, even with the same number of requirements. After we reformatted all game prompts into similar formats with the optimizer, the correlation between requirement and LLM output for Connect4 increased from $\rho=0.02$ to $\rho=0.3$, partially supporting our hypothesis. In contrast, Tic-Tac-Toe did not exhibit such sensitivity, likely due to a broader variety of rules in LLM’s training data. GPTs tasks did not have this problem since both the prompt and implementation are in natural language.
We further analyzed how different types of requirement defects impact LLM output quality. We found that the numbers of omission errors (incomplete requirements) had a stronger negative impact on LLM Output Quality Score ($\rho=-0.4$) compared to commission errors (inaccurate requirements) ($\rho=0$). This suggests that while LLMs can correct inaccuracies in requirements, they struggle to fill in missing information — further highlighting the importance of training humans to express all their necessary requirements.
Our findings complement existing existing instruction following research (Zhou et al., 2023; Jiang et al., 2023), which tend rely on synthetic prompts with perfectly written requirements of a fixed set of common constraint types (Section 2.2). Our results highlight the need for further investigation into how imperfect user requirements — both in content and format — interact with inherent LLM biases.
@@
Finally, we evaluated how the optimizer affected prompt quality, experimenting on a single optimizer to explore feasibility, without generalizing to all optimizers (Section 7.1).
We first noticed that the Prompt Maker optimizer effectively enhanced user prompts by adding elements such as role-playing and chain-of-thought, making prompts more structured and fluent, as shown in Figure 2. Besides formatting, the optimizer also made direct modifications to requirements; on average, it added 0.5 missing requirements and corrected 0.3 inaccurate requirements, reducing omission errors from $4.4$ to $3.9$ and commission errors from $0.6$ to $0.3$ per participant across all prompts. These changes significantly improved overall test scores, as shown by a paired t-test for all users’ prompts before and after optimization ($p<0.001$).
It seemed like there exist some requirements that were LLM-hard but not optimizer-hard. For example, GPT-4o would not implement ‘‘Display messages for a tie game’’ unless explicitly prompted, but the optimizer would automatically add this requirement to user prompts. However, highly customized requirements, such as ‘‘Cross out the winning cells’’, were never automatically added (0 out of 51) without explicit user input.
Importantly, the optimizer could not close the performance gap between the ROPE and PE group. We computed the gains of PE participants after using the optimizer ($8.8\%$ = optimized prompts’ post-test scores -- original prompts’ pre-test scores). Using a two-sample t-test, the ROPE’s 19.8% pre- to post-test gains remained significantly higher ($p=0.01$) than PE group’s optimized gains ($8.8\%$). This supports the finding that missing, highly customized requirements are unlikely to be added automatically.
@@
These issues highlight gaps in user-expressed requirements and LLM-received requirements, suggesting that future training should address the risks of misinterpretation by either the LLM or the optimizer (Section 7.2). On a more positive note, such misinterpretations could provide useful feedback in the future, alerting users to potential misalignment between their requirements and the LLM’s understanding.
@@ 7. Discussion
In this work, we introduce Requirement-Oriented Prompt Engineering (ROPE), a human-AI collaborative prompting paradigm where humans focus on effective requirements specification. Our evaluation shows that requirement-focused training significantly enhances novices’ ability to extract and articulate requirements in prompts, leading to more goal-aligned LLM outputs; we also notice mismatch between humans’ under-specified requirements and LLMs’ misinterpretations is a key communication barrier (Section 6.2). While our work made an important step toward ROPE, several limitations remain, and there are many interesting questions to explore. Here, we reflect on possible immediate extensions on requirement-focused training, as well as the longer-term evolution of the ROPE paradigm.
@@ 7.1. Limitations and Next Steps on Requirement-Focused Training
Our training program significantly improved participants’ ability to extract and articulate requirements, though the skill remains challenging. Post-training assessments showed an average score of 40%, with the top performer increasing from 24% to 63%. To further improve the effectiveness of requirement-focued training, we propose several next steps.
@@
In addition to extending length, broadening the scope of the training will help improve the generalizabilty. Our study showed that the skills gained transferred to new GPTs and game development tasks. However, an open question remains: how well do these skills apply to other tasks like data analysis or creative content generation? Fortunately, our ROPE interface is adaptable to different tasks, as training can be generated based on provided reference requirements and LLM outputs. We plan to open-source ROPE and encourage its use for customized tasks across domains.
Besides length, our training can also be broadened for better generalzability. Our study demonstrated that the requirement specification skills users acquired through training generalized to new GPTs and Game development tasks. However, an open question remains: how broadly do these requirement skills transferred to other natural language programming tasks (e.g., data analysis or creative content generation)? Fortunately, our ROPE interface is adaptable to different tasks, as the training and feedback loop can be easily generated based on provided reference requirements and LLM outputs. We plan to open-source ROPE and encourage its use for customized tasks across domains. Moreover, beyond human-AI interactions, ROPE skills — such as requirement extraction, iterative refinement — can be broadly applicable to fields like software engineering and design. Future research should explore further transfer tasks in assessments, and also how ROPE skills may enhance communication, problem decomposition, and computational thinking across various fields.
Another dimension of generalizability is towards users’ different prompting behaviors. Our study focused on reverse engineering fixed requirements, which allowed for controlled feedback but did not fully capture real-world prompting scenarios: users often engage in one-off interactions (e.g., direct question-answering) or iterative and open-ended prompting (e.g., chatting with LLMs on a rough idea), etc. (Shankar et al., 2024a; Kim et al., 2023). Future studies should investigate more self-directed tasks, compare different types of prompting, and tailor training accordingly. While providing feedback without clear requirements may be challenging, pre-generated materials on common novice approaches could offer useful scaffolding.
One primary bottleneck for broadening the scope is our lack of automatic assessment methods for the LLM-hard tasks. Our time-consuming manual grading effort limited our ability to conduct more experiments by e.g., grading multiple LLM outputs per user prompt for stability, experimenting with multiple optimizers, etc. Future work should explore automatic assessments and their trade-offs. For example, LLM-as-a judge paired with few-shot prompting might provide reasonable estimations (Zhao et al., 2021), and might be suitable for tasks that contain substaintial numbers of requirements where wrong gradings on a single requirement dot not drastically affect the assessment results.

### v2

@@ 5. User Study Design
To capture the requirement-focused training gains, we utilized a pre-test and post-test approach, a standard instruction evaluation method (Rogaten et al., 2019), comparing participants’ performance on isomorphic tasks before and after ROPE training (within-subject). To compare requirement-focused training with standard training, we utilized a randomized-control design (between-subject). We randomly assigned participants to either the experimental condition (ROPE group), where they receive requirement-focused training in our interface, or the control condition (PE group), where they receive a standard prompt engineering tutorial and then self-practice using ChatGPT.
Our user study — for either the ROPE or the PE group — is 1.5 hours long. In the study, participants first completed a two-minute pre-survey, providing demographic information and rating their prompting experiences (e.g., confidence and self-perceived productivity using ChatGPT) with 7-level Likert Scale questions. Then, they started a pre-test (20 minutes) to write prompts to replicate key features on existing programs (Table 1) — one Game (Tic-Tac-Toe or Connect4) and one GPTs (Outline Assistant or Trip Advisor). They proceeded to the 40-minute training session (different interfaces depending on experiment group) on two tasks (Tetris and Email Proofreader), followed by a post-test that contained different Game and GPTs tasks (20 minutes).141414 The pre- and post- assessments each contained one Game task and one GPTs task (Section 4.1), counterbalanced to mitigate problem sequencing bias or different task difficulties. We found that the two versions of the tests have comparable difficulties from pre-test results (an average overall score of 21.0% and 25.6% for each version). To further capture participants’ iterative prompting behaviors, for the post-test GPTs task, participants were also asked to feed their prompt into ChatGPT and iterate their original prompt by observing the ChatGPT outputs. They completed the study with another two-minute post-survey, providing feedback with open-ended or Likert Scale questions on their experience and perceptions during the training. Participants were compensated with a \$20 Gift Card. Please refer to supplemental materials for our study materials.
@@
To understand participants’ perceived experiences, prompting behaviors, and learning outcomes, we collected participants’ prompt responses, survey answers, time on task, as well as interaction log data during the training phase. Afterward, we collected the LLM outputs by feeding participant prompts into OpenAI GPT-4o. [^16] We used a temperature of 0 to generate deterministic outputs, please refer to our supplemental materials for more details. Using the evaluation method and rubrics we developed for each of the assessment tasks Section 4.2, we graded the pre- and post- tests using users’ original prompts for all tasks from the test, on a scale of 0% to 100%. We calculated the Overall Scores as the average of the Requirement Quality and LLM Output Quality scores.
To understand the effect of an optimizer on a prompt (as discussed in Section 2.2), we further use the Prompt Maker, [^17] https://chatgpt.com/g/g-hhh4w3eov-prompt-maker: from a simple prompt to an optimized prompt. a popular optimizer with 50k usages in GPT Store, to refine participants’ original prompts into optimized prompts. This optimizer is easily applicable to any kind of prompts (Figure 2 shows an example of its output), though it is a rather preliminary version, as it rephrases user prompts in a rule-based manner without criticizing the LLM outputs of the original prompt. We evaluated the optimized prompts in the same way as the human prompts (Section 4.2), including generating and grading corresponding LLM outputs.
@@
We recruited users with limited or no NLP background from different institutions in the United States. We conducted studies with 32 participants randomly assigned to the ROPE or PE group, and we removed two participants who disengaged during the study from the analysis. In total, we have 30 participants (S1-30, $n=15$ for each condition) — 19 female, 10 male, 1 agender, and 18 non-native English speakers (9 in each condition), with an average age of 26. Our participants have a diverse range of backgrounds including HCI, education technology, communication, graphic design, mechanical engineering, information systems, literature, psychology, and economics. We asked the participants to rate their familiarity with LLMs and complex prompts in the pre-survey, and the average rating is 3 out of 7.
@@ 6.1. Learning Gain: Requirement-focused Training Helps Students Instruct LLMs
Using a two-tailed paired t-test, [^18] We did a linear mixed effects regression to account for the correlation across participants and task (both as random effects) of the repeated measures for the pre- and post-test times and for the requirement and LLM output scores. We found the same pattern of significant results as revealed by the t-tests. we found that the overall scores for participants in the ROPE group significantly improved by 19.0% [^19] Requirement Quality score improved by 25.4% and LLM Output Quality score improved by 12.7%, both significant ($p<0.05$). from pre- to post-test ($p<0.001$, from $21.9\%\pm 14.6\%$ to $40.9\%\pm 18.4\%$), an almost two-fold increase. This demonstrated that requirement-focused training enabled participants to instruct LLMs more effectively, and we achieved our desired goal: to concentrate users’ attention on only requirement iterations during the training, but make sure users can still apply their learning in overall prompt writing. Analyzing requirement defects in ROPE participants’ prompts, we observed a noticeable decrease in omission errors (from 5.6 to 3.2 per participant), but a slight increase in commission errors (from 0.5 to 0.7 per participant). This indicated that participants became more aware of requirements, but need additional training for requirements clarity and accuracy.
ROPE participants’ self-reflection highlighted that they understand the value of emphasizing and iterating on clear and complete requirements, which may contribute to their improvement. For example, 7 out of 15 PE group and 13 out of 15 ROPE group participants mentioned “be specific” in their answers to prompting strategies in post-survey. When asked about what they learned during the instruction, four PE participants noted to “give exact instruction,” as described by S23, “we should give as much information as we can to ChatGPT”. Meanwhile, some ROPE participants developed a more in-depth understanding of articulating requirements in prompts. For example, six ROPE participants explicitly noted the connection between requirements and prompts. S30 described requirements as a way “to be more clear and not confuse the system by giving too much unnecessary information”, and S8 noted that a requirement-focused approach helped them organize their thoughts: “I learned to organize my requirements logically so we can easily revise and improve them”. S10 neatly described their shift towards requirement-focused prompting strategy: “sometimes when I write prompts, [I find] steps are hard to be clearly divided, or I didn’t consider to divide them that detailed. However, it’s important to do so, as it appears to give LLM more direct and clear instruction. When the steps are divided, it’s easier to see the missing details in my original prompts too”.
@@
These prompt also revealed learning gain differences. A two-sample t-test showed that the ROPE achieved significantly higher learning gains compared to the PE group ($p<0.001$). In fact, while ROPE significantly improved (19.0% as mentioned above), we did not observe much gain in PE group’s pre-to-post score (0.7%). The contrast suggested that novices indeed could not acquire requirement articulation skills through standard prompting training alone.
@@
Requirement articulation was challenging for all participants, as shown by the low pre-test scores (21.9% for the ROPE and 24.8% for the PE). This poor starting point likely hindered the PE’s ability to receive useful feedback during self-practice, as their requirements were too weak for the LLM to generate “good enough” model output with clear indicators on what to improve. Consequently, PE participants were trapped and frustrated by LLM unpredictability — S6 noted that “ChatGPT is a very malleable tool and can change responses pretty drastically depending on the prompt”, and S7 commented that “if ChatGPT doesn’t get it, I might not be patient”.
In contrast, our feedback loop for ROPE (Section 4.3) provided tightly controlled feedback and program output based on the requirements users specified, regardless of grammatical issues or natural language variances. S8 highlighted that: “It’s easy to understand (very logical) and can see how the changes and revises influence the system immediately”. This deterministic feedback reinforced ROPE participants’ belief in the importance of clear requirements, leading to them writing better initial prompts with more requirements in the post-test. We hypothesize that these better initial prompts might also become more suitable for further iteration. Although still imperfect, ROPE participants’ prompts might allow the LLM to generate partially correct outputs that contain enough signal towards future improvements.
@@ 6.2. In-depth Analysis: The Validity of ROPE Paradigm
Table 3. Spearman correlations ($\rho$) between LLM Output Quality and Requirement Quality scores by tasks
Task Type Task $\rho$
GPTs Task TripAdvisor 0.79
OutlineAssistant 0.75
Game Task TicTacToe 0.90
Connect4 0.08
Overall 0.71
Going beyond learning gains, we dive deeper into the connections between requirements in prompts and LLM output quality, as well as the role of optimizers.
@@
To examine whether more correct and complete requirements led to better LLM outputs, we calculated the correlation between the two components of our Overall Scores: Requirement Quality and LLM Output Quality scores. We found a strong positive correlation, with a Spearman’s correlation coefficient of $\rho=0.71$ (Schober et al., 2018). However, task-specific analysis revealed nuances. As shown in Table 3, while most tasks achieved $\rho\geq 0.7$, Connect4 showed no correlation. Upon further inspection, we found that interestingly, the LLM tends to produce two implementation types for Connect4: command line and pygame. Prompts that include interactive or visual requirements such as ‘‘Highlight the hovered column’’ often led to pygame implementations. If we split the Connect4 correlations by the LLM output implementation type, we get reasonable numbers for both cases: $\rho=0.89$ (command line) and $\rho=0.58$ (pygame), suggesting that implementations may respond to requirements differently. For example, the requirement ‘‘Display player’s turn message’’ was always implemented in command line Connect4, even if omitted in prompts. However, only 1 pygame Connect4 implemented the same requirement, out of 6 participants who correctly described it. This suggests that requirements can be interdependent, with LLM-hardness varying by implementations.
We further analyzed how different types of requirement defects impact LLM output quality. We found that the number of omission errors (incomplete requirements) had a stronger negative impact on LLM Output Quality Score ($\rho=-0.49$) compared to commission errors (inaccurate requirements) ($\rho=0.03$). This suggests that while LLMs can correct inaccuracies in requirements, they struggle to fill in missing information — further highlighting the importance of training humans to express all their necessary requirements.
Our findings complement existing instruction following research (Zhou et al., 2023; Jiang et al., 2024), which tend to rely on synthetic prompts with perfectly written requirements of a fixed set of common constraint types (Section 2.2). Our results highlight the need for further investigation into how imperfect user requirements interact with inherent LLM biases.
@@
Finally, we evaluated how the optimizer affected prompt quality, experimenting on a single optimizer to explore the feasibility, without generalizing to all optimizers (Section 7.1).
We first noticed that the Prompt Maker optimizer effectively enhanced user prompts by adding elements such as role-playing and chain-of-thought, making prompts more structured and fluent, as shown in Figure 2. Besides formatting, the optimizer also made direct modifications to requirements; on average, it added 0.5 missing requirements and corrected 0.3 inaccurate requirements, reducing omission errors from $4.4$ to $3.9$ and commission errors from $0.6$ to $0.3$ per participant across all prompts. These changes significantly improved overall scores, as shown by a paired t-test for all users’ prompts before and after optimization ($p<0.001$).
It seemed like there existed some requirements that were LLM-hard but not optimizer-hard. For example, GPT-4o would not implement ‘‘Display messages for a tie game’’ unless explicitly prompted, but the optimizer would automatically add this requirement to user prompts. However, highly customized requirements, such as ‘‘Cross out the winning cells’’, were never automatically added (0 out of 51) without explicit user input.
Importantly, the optimizer could not close the performance gap between the ROPE and PE group. We computed the gains of PE participants after using the optimizer ($7.3\%$ = optimized prompts’ post-test scores -- original prompts’ pre-test scores). Using a two-sample t-test, the ROPE’s 19.0% pre- to post-test gains remained significantly higher ($p=0.01$) than PE group’s optimized gains ($7.3\%$). This supports the finding that missing, highly customized requirements are unlikely to be added automatically.
@@
These issues highlight potential gaps in user-expressed requirements and LLM-interpreted requirements, suggesting that future training should address the risks of misinterpretation by either the LLM or the optimizer (Section 7.2). On a more positive note, such misinterpretations could provide useful feedback in the future, alerting users to potential misalignment between their requirements and the LLM’s understanding.
@@ 7. Discussion
In this work, we introduce Requirement-Oriented Prompt Engineering (ROPE), a human-AI collaborative prompting paradigm where humans focus on effective requirements specification. Our evaluation shows that requirement-focused training significantly enhances novices’ ability to extract and articulate requirements in prompts, leading to more goal-aligned LLM outputs; we also notice a mismatch between humans’ under-specified requirements and LLMs’ misinterpretations is a key communication barrier (Section 6.2). While our work made an important step toward ROPE, several limitations remain, and there are many interesting questions to explore. Here, we reflect on possible immediate extensions on requirement-focused training, as well as the longer-term evolution of the ROPE paradigm.
@@ 7.1. Limitations and Next Steps on Requirement-Focused Training
Our training program significantly improved participants’ ability to extract and articulate requirements, though the skill remains challenging. Post-training assessments showed an average score of 41% (max=63%), and the top ROPE learner gained 43% in scores from pre- to post-test. To further improve the effectiveness of requirement-focused training, we propose several next steps.
@@
In addition to extending length, our training can also be broadened for better generalizability. Our study demonstrated that the requirement specification skills users acquired through training generalized to new GPTs and Game development tasks. However, an open question remains: how broadly do these requirement skills transfer to other natural language programming tasks (e.g., data analysis or creative content generation)? In preparation for generalizability, our ROPE interface is adaptable to different tasks, as the training and feedback loop can be easily generated based on provided reference requirements and LLM outputs. We open-source ROPE system and encourage its use for customized tasks across domains. Moreover, beyond human-AI interactions, ROPE skills — such as requirement extraction and iterative refinement — can be broadly applicable to fields like software engineering and design. Future research should explore further transfer tasks in assessments, and also how ROPE skills may enhance communication, problem decomposition, and computational thinking across various fields.
Another dimension of generalizability is towards users’ different prompting behaviors. Our study focused on extracting fixed requirements, which allowed for controlled feedback and simulated the scenarios when interaction examples are available. However, this setup did not fully capture all real-world prompting scenarios: users often engage in one-off interactions (e.g., direct question-answering) or iterative and open-ended prompting (e.g., chatting with LLMs on a rough idea), etc. (Shankar et al., 2024a; Kim et al., 2024). Future studies should investigate more self-directed tasks, compare different types of prompting, and tailor training accordingly. While providing feedback without clear requirements may be challenging, pre-generated materials on common novice approaches could offer useful scaffolding.
One primary bottleneck for broadening the scope is our lack of automatic assessment methods for the LLM-hard tasks. Our time-consuming manual grading effort limited our ability to conduct more experiments by e.g., grading multiple LLM outputs using different temperatures per user prompt, experimenting with multiple optimizers, etc. Future work should explore automatic assessments and their trade-offs. For example, LLM-as-a judge paired with few-shot prompting might provide reasonable estimations (Zhao et al., 2021) and might be suitable for tasks that contain substantial numbers of requirements where wrong grading on a single requirement does not drastically affect the assessment results.